# On Modeling and Solving the Boltzmann Equation

### v1

@@ 7.2 Kinetic Models and the G-problem
In solving a wide class of problems in the rarefied gas dynamics listed in this Section, algebraic manipulations led us to the application of the ADO method to the solution of an auxiliary problem we usually referred to as the “G-problem” that we write here in the form
$\xi{\partial\over\partial\tau}G(\tau,\xi)+G(\tau,\xi)=\int_{-\gamma}^{\gamma}G(\tau,\xi^{\prime})\,d\xi^{\prime}\,$, (119)
@@
for $\xi\in(0,\gamma)$. Here, $\Psi(\xi)$ is called the characteristic function, which assumes different definitions depending on the problem. We note that for some applications (Barichello and Siewert, 1999b; Barichello et al., 2001), $\gamma$ is such that the integration interval is defined as $(-\infty,\infty)$. The central aspect we want to point out is that the challenging solution of a “G-problem” type was the motivation for proposing the ADO method (Barichello and Siewert, 1999b). The first paper with the derivation of the ADO method. Even more, the versatility of the ADO formulation in solving this more general type of problem (than the problem given by Eqs. (7) and (10)) distinguishes the ADO method from other discrete ordinate methods usually associated only with the treatment of problems in the area of neutron transport and radiative transfer. It is worth to note that, the research on the kinetic models led to the solution of the LBE (Siewert, 2004a; Garcia and Siewert, 2010) through the application of the ADO method to the scalar “G problem”, Eq. (119), and its vector versions.
@@ 8 Inverse Problems
The focus of this work was the development and application of solutions, using the ADO method, to solve direct problems. However, there are many inverse problems of interest in transport theory. For instance, we have investigated the reconstruction of radiation sources, which are fundamental to nuclear safety, in recent years (Pazinatto et al., 2020; Pazinatto and Barichello, 2021). Currently, we are studying the problem of reconstructing the absorption and scattering coefficients, which is essential to optical tomography (Groff et al., 2025).
@@ 9 Concluding Remarks and Ongoing Research
In this work, a brief overview of studies and advances related to the solution of the discrete ordinates approximation of the linear Boltzmann equation in one- and two-dimensional spatial dimensions was presented. The focus was on the application areas of neutrons, photons, and rarefied gas dynamics. In particular, the successful applicability of the Analytical Discrete Ordinates (ADO) method to a series of particle transport problems was emphasized through the reference of selected relevant publications. A broad literature review in the area is not within the scope of this work, but rather to draw attention to research in this important area, carried out in Latin America and more specifically in Brazil.
Despite the advances and contributions we seek to highlight, several important issues still deserve attention for the development of the method and the subject. In particular, domain decomposition methods are being investigated to improve the large-scale linear system issue that is part of the formulation. Another area of study, still open, is the analysis of asymptotic error due to angular discretization—more specifically, the joint analysis of spatial-angular discretizations. Finally, we intend to use the formulation with exact representation of the phase function in the problems of reconstruction of absorption and scattering parameters, of interest in optical tomography, treated by the ADO method, so far, only in one-dimensional geometry.
Acknowledgments The author thanks to CNPq of Brazil for partial support to this and other works.

### v2

@@ 7.2 Kinetic Models and the G-problem
In solving a wide class of problems in the rarefied gas dynamics listed in Section 7, algebraic manipulations led us to the application of the ADO method to the solution of an auxiliary problem we usually referred to as the “G-problem” that we write here in the form
$\xi\frac{\partial}{\partial\tau}G(\tau,\xi)+G(\tau,\xi)=\int_{-\gamma}^{\gamma}G(\tau,\xi^{\prime})\,d\xi^{\prime}\,$, (119)
@@
for $\xi\in(0,\gamma)$. Here, $\Psi(\xi)$ is called the characteristic function, which assumes different definitions depending on the problem. We note that for some applications [17, 7], $\gamma$ is such that the integration interval is defined as $(-\infty,\infty)$. The central aspect we want to point out is that the challenging solution of a “G-problem” type was the motivation for proposing the ADO method [17]. The first paper with the derivation of the ADO method. Even more, the versatility of the ADO formulation in solving this more general type of problem (than the problem given by Eqs. (7) and (10)) distinguishes the ADO method from other discrete ordinate methods usually associated only with the treatment of problems in the area of neutron transport and radiative transfer. It is worth to note that, the research on the kinetic models led to the solution of the LBE [60, 37] through the application of the ADO method to the scalar “G problem”, Eq. (119), and its vector versions.
@@ 8 Inverse Problems
The focus of this work was the development and application of solutions, using the ADO method, to solve direct problems. However, there are many inverse problems of interest in transport theory. For instance, we have investigated in recent years the reconstruction of radiation sources, which are fundamental to nuclear safety [51, 50]. Currently, we are studying the problem of reconstructing the absorption and scattering coefficients, which is essential to optical tomography [38].
@@ 9 Concluding Remarks and Ongoing Research
In this work, a brief overview of studies and advances in the solution of the discrete ordinates approximation of the linear Boltzmann equation in one and two spatial dimensions was presented. The focus was on the application areas of neutrons, photons, and rarefied gas dynamics. In particular, the successful applicability of the Analytical Discrete Ordinates (ADO) method to a series of particle transport problems was emphasized through the reference of selected relevant publications. A broad literature review in the area is not within the scope of this work, but rather to draw attention to research in this important area, conducted in Latin America and more specifically in Brazil.
Despite the advances and contributions we have sought to highlight, several important issues still deserve attention for the development of the method and the subject. In particular, domain decomposition methods are being investigated to address the large-scale linear system issue that is inherent to the formulation. Another area of study, still open, is the analysis of asymptotic error due to angular discretization — more specifically, the joint analysis of spatial-angular discretizations. Finally, we intend to use the formulation with exact representation of the phase function in the problems of reconstruction of absorption and scattering parameters, of interest in optical tomography, treated by the ADO method, so far, only in one-dimensional geometry.